OpenStudy (anonymous):

Divide. 3 1/9 divided by 2 1/3 = 3 3/4 divided by 4 2/7 = 4 1/2 divided by 3/8 = 8 3/4 divided by 1/3 = 3 5/6 divided by 2 2/3 = 51/2 divided by 3 1/8 =

OpenStudy (anonymous):

Medals for ALL answers!

OpenStudy (anonymous):

\(2\tfrac{1}{3}=\frac{7}{3}\) \[\frac{1}{9}\div\frac{7}{3}=\frac{1}{9}\times \frac{3}{7}=\frac{1}{21}\] others area similar